# Digital currency hardware wallets and the essence of money

### 28 Feb 2026 v4

@@ Acknowledgements
The author thanks Professor Tomaso Aste for his continued support for his ongoing work on digital currencies and digital payment systems. The author thanks Christian Grothoff and Yuri Biondi for their generous and insightful feedback on earlier drafts of this article. The author also thanks his many colleagues in the international standards development community for a series of vigorous arguments that were instrumental to reaching the crux of this most contentious debate. The author also acknowledges the Future of Money Initiative at University College London, the Systemic Risk Centre at the London School of Economics, EPSRC, and the PETRAS Research Centre EP/S035362/1 for the FIRE Project.

### 12 Mar 2026 v5

@@ Acknowledgements
The author thanks Professor Tomaso Aste for his continued support for our ongoing work on digital currencies and digital payment systems. The author thanks Christian Grothoff and Yuri Biondi for their generous and insightful feedback on earlier drafts of this article. The author also thanks his many colleagues in the international standards development community for a series of vigorous arguments that were instrumental to reaching the crux of this most contentious debate. The author also acknowledges the Future of Money Initiative at University College London, the Systemic Risk Centre at the London School of Economics, EPSRC, and the PETRAS Research Centre EP/S035362/1 for the FIRE Project.